Journalist Report – March 5th

Crew 206 – 03/05/2019

Sol 9

Author: Benjamin Auzou, Journalist

“Martian Wild Life”

Second day of our second week here in the MDRS, also the return of the routine ! We began the day with our sport session and yesterday and today, we all broke our records of the previous week. We are on the way to the roof of the Olympus (Mons).

After climbing the slopes of the western area of the Hab, the EVA team of the day walked along Hab Ridge. Aurélien, Norbert and Jérémy were the last crewmembers that never went here, after Cerise, Gaspard and I explored this zone last week. On their walk around the area, they spotted three strange animals. Under the surprise of that discovery, they didn’t manage to identify the species.
Jérémy took advantage of the view up there to shoot the whole landscape in order to create a panoramic view from the ridge. Such places are perfect for photo passionates like him, the colors and the relief of the day and the clear and pure sky of the night are beautiful to capture. The result is wonderful, a large picture of the Hab and the landscape around it : 170 MB of red dust and blue sky.

After three days of hard time trying to finish the Mars puzzle, Cerise, Gaspard and I finally managed to assemble the thousand pieces and we wrote our crew name with Crews 198, 203 and 205 at the back of the puzzle. When we added the last piece to it, an exclamation of joy echoed in the Hab. Crew 207 and following crews, we challenge you to assemble the puzzle during your rotation !

In the afternoon, the Green-Hab was musical ! Indeed Norbert continued his experiment on the influence of music and frequencies on plants resistance under radiations. We’re waiting for his first results that will certainly fall in the last days of the mission.

This night will mark the mid-mission milestone for us. As we turn in the second part of this extraordinary adventure, the crew is still holding strong together, and seeing the end of this journey coming both excites and frightens us.

Operations Report – March 5th

Crew 206 Operations Report 05 – MAR – 2019

SOL: 9

Name of person filing report: Gaspard Thieulin

Non-nominal systems: Generator

Notes on non-nominal systems: Radiator hoses and
alternator need replacement on generator. A contractor has
been arranged to make these repairs within a few days.

Generator (hours run): 14,5 hours

Solar— SOC % – Turned off Gen : 100%, Turned on Gen : 81%

SOC tracking :
SOL 9 (MAR 05th) :
9h30: 100%
12h20: 92%
13h30: 90%
16h00: 86%
18h00: 82%

Diesel Reading – 70%

Propane Reading – 80%

Ethanol Free Gasoline – 7 gallons

Water (Auxillary tank) – Not in use

Water (Static tank) – 425 gallons

Water differential (static tank) – 13 gallons

Gallons transferred: Not Applicable

Water in GreenHab – 245 gallons

Static to Loft Pump used – Yes

Water Meter: 0141666,0 gallons

Toilet tank emptied: No

Deimos rover used: No, still not functional
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Not Applicable
Beginning charge: Not Applicable
Ending charge: Not Applicable
Currently charging: Not Applicable

Opportunity rover used: Not used
Hours: 53,0 hours
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Curiosity rover used: Used
Hours: 77,4 hours
Beginning charge: 100%
Ending charge: 98%
Currently charging: Yes

Spirit rover used: Used
Hours: 78,4 hours
Beginning charge: 100%
Ending charge: 94%
Currently charging: Yes

Notes on rovers: Nothing to report.

ATV’s Used: N/A

Reason for use: Not used

Oil Added? N/A

ATV Fuel Used: N/A

# Hours the ATVs were Used today: None

Notes on ATVs: None

Hab Car used and why, where? Not used.

Staff Crew Car used and why, where? Off Campus

General notes and comments: Nothing to report.

Summary of internet: “Old” router was plugged in in place of the other one. No notable difference. A lot of data sinking today, the only new device connected to the Internet is the camera system.

Summary of suits and radios: Nothing to report.

Summary of Hab operations: Nothing to report.

Summary of Green Hab operations: Nothing to report.

Summary of Science Dome operations: Nothing to report.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Main airlock’s lock is broken: the door part doesn’t hold on the door anymore. I’ll look for one in the hab, but we could use a new spare one at least.

EVA Report – March 5th

Crew 206 EVA Report 05 Mar 2019
EVA #07

Author : Jérémy Auclair

Purpose of EVA: The EVA has two purposes :

– Maintenance of the LOAC, installation of a rain protection

– Exploration and photography of the Hab ridge

Start time: 9:15 am

End time: 11:15 am

Narrative: We were happy to discover that the ground was very dry this morning. At 09:20 am, after having done the checks around the Hab, we took Spirit and Curiosity to go to the LOAC with some equipment. We installed a very simple rain protection for the LOAC, it’s probably not as effective as I’d like, but it’s better than nothing for now. At 09:40 am we parked Spirit and Curiosity and started walking north to climb up the ridge. We took a little less than 15 minutes to arrive to the top. The view is breathtaking, even after a year. Once on top we walked south for 25 minutes, taking picture along the way. It’s a very interesting spot because there are very different landscapes that join to this point.
After the 25 minutes walk, we turned around to find a way to go back down, we walked for 10 minutes. We walked on a south oriented light slope, for maximum security. The way down was easy, the ground was dry, everything went fine. We arrived at the Hab at 11:10 am and entered the main airlock.

Destination: LOAC and Weather Station location, Hab Ridge

Coordinates: 518500/425700 (LOAC and Weather Station), 517900/4251200 (Hab Ridge Road /Sagan Street)

EVA Participants: Norbert Pouzin (GHO), Aurélien Mure (XO), Jeremy Auclair (CMD, EVA leader)

Roads and routs per MDRS Map: Cow Dung Road

Mode of travel: Driving and Walking

Greenhab Report – March 5th

Crew 206 Green Hab Report – 05-MAR-2019
GreenHab Officer: Norbert Pouzin

Environmental control: Heating and Cooling w/ ambient air (6 hrs)

Shade Cloth (40%)

Average temperatures:

Low: 17.0°C

High: 25.4°C

Hours of supplemental light: 0 Hours

Daily water usage for crops: 4.38 gallons

Water in Blue Tank: 245.2 gallons

Crops watered at 0830 hours (1.73 gallons) and 1815 hours (2.65 gallons):


1) 08.30 AM 18.8° C / 30 %
2) 12.00 AM 22.5° C / 24 %
3) 15.00 PM 19.4° C / 30 %
4) 18.15 PM 18.9° C / 38 %

Changes to crops: Cress is sprouting.

Today wasn’t as hot as yesterday. I turned off the fan earlier, and only opened the door until 15.00.
I continued my experiment in the greenhab this afternoon. Everything went well.

Harvest: None.

Support/supplies needed: None.

Sol Summary – March 5th

Crew 206 March 5th 2019
Sol 9

Summary Title : Sleepy Commanders

Author : Jeremy Auclair

Mission Status : All nominal

Sol Activity Summary : Yesterday night we went to the observatory to take a look at the sky and take some pictures. There was almost no clouds, and the air was pretty clear. I had an issue with my camera lens, so I only took a couple nice pictures. As usual everyone got out of bed at 07:00 am for the daily workout. It’s still as hard, even if we’re getting used to it. After that we ate breakfast and prepared for the EVA. We went to attach a rain protection to the LOAC. It’s very simple, probably not as effective as I’d wish, but it’s good enough for now. After that we went up to the Hab ridge. The view is still as impressive and beautiful, even after a year. I took some great pictures. Cerise felt sick this morning, so she didn’t go on the EVA, and took advantage of this time to finish the Mars puzzle (unfortunately there is 3 missing pieces). After lunch, 5 of us took a long nap. It didn’t help me that much, I’m feeling sleepy since then and Aurélien, our XO, feels also quite sleepy. I haven’t made some progress on TELEOP today, our extra charger is supposed to arrive tomorrow. At least I have a set-up that works, I’m gathering a part of the data I planned on gathering. After that is was a quiet day.

Look ahead plan : We are now planning all our filming sessions for the mission video, it will take some time and some work, but it should look good !

Anomalies in work : None

Weather : Sunny in the morning, Cloudy in the afternoon

Crew physical status : Fine

EVA : LOAC maintenance and Hab Ridge exploration

Reports to be filed : GreenHab, Journalist, Operation, EVA Report, EVA Request

Support Requested : None

Copyright © The Mars Society. All rights reserved. | Main Site